2x−2, KL=x-9,KL=x−9, and JL=2x+8,JL=2x+8, determine the numerical length of \overline{KL}.
KL
.
1
Expert's answer
2020-11-19T16:29:22-0500

KL = x-9

JL = 2x+8

I assume JK = 2x-2(probably a typographical error made "JK" to be cut out of the question)


KL segment + JK segment = JL segment

x-9 + 2x-2 = 2x+8

3x-11 = 2x + 8

x = 19

Therefore:

The numerical length of KL = x - 9

= 19 - 9

=10